# DeepProtein: Deep Learning Library and Benchmark for Protein Sequence Learning

### v1

@@ Abstract
In recent years, deep learning has revolutionized the field of protein science, enabling advancements in predicting protein properties, structural folding and interactions. This paper presents DeepProtein, a comprehensive and user-friendly deep learning library specifically designed for protein-related tasks. DeepProtein integrates a couple of state-of-the-art neural network architectures, which include convolutional neural network (CNN), recurrent neural network (RNN), transformer, graph neural network (GNN), and graph transformer (GT). It provides user-friendly interfaces, facilitating domain researchers in applying deep learning techniques to protein data. Also, we curate a benchmark that evaluates these neural architectures on a variety of protein tasks, including protein function prediction, protein localization prediction, and protein-protein interaction prediction, showcasing its superior performance and scalability. Additionally, we provide detailed documentation and tutorials to promote accessibility and encourage reproducible research. This library is extended from a well-known drug discovery library, DeepPurpose and publicly available at https://github.com/jiaqingxie/DeepProtein/tree/main.
@@ 1 Introduction
Understanding the representation of proteomics is vital in developing traditional biological and medical progress (Wu et al., 2022b; Fu et al., 2024), multi-omics genomics (Wu et al., 2022a; Chen et al., 2021), and curing human diseases (Chen et al., 2024c, b). Being the working house of the cell, it provides lots of functions that support human’s daily life, such as catalyzing biochemical reactions that happen in the body as a role of enzymes and providing helpful immune responses against detrimental substance that acts as immunoglobulin (Wu et al., 2024). Under the necessity of analyzing those useful proteins, several related protein databases are available to researchers (Berman et al., 2000; Bairoch and Apweiler, 2000; Consortium, 2015; Pontén et al., 2008). Apart from the 2D database, some recent 3D Protein Database used AlphaFold 2.0 (Jumper et al., 2021) is important to better assist in learning those representations in 3d-dimensional space. The success of AlphaFold 2.0 has sparked a significant increase in interest in using machine learning techniques for protein learning tasks, of which the goal is to improve our understanding of proteins’ biochemical mechanisms.
Deep learning has shown its power in protein learning tasks, including protein-protein interaction (Gainza et al., 2020), protein folding (Jumper et al., 2021; Lu, 2022; Panou and Reczko, 2020; Chen et al., 2016), protein-ligand interaction (Li et al., 2021b; Xia et al., 2023), and protein function and property prediction (Gligorijević et al., 2021; Sevgen et al., 2023). Convolutional neural network (CNN) (Shanehsazzadeh et al., 2020) and TAPE Transformer (Rao et al., 2019) on proteins are models with regard to protein sequence-to-sequence (Seq2Seq) learning, including the pretrained transformer such as ProtBert (Brandes et al., 2022). With the help of 3D information intrinsics, graph neural networks (GNN) are applied, especially some pretrained GNN models (Jing et al., 2020; Zhang et al., 2022), and graph transformers (Yuan et al., 2022; Gu et al., 2023). While transformer claims to be the state-of-the-art in the previous benchmark (Xu et al., 2022), the analysis of potential powerful deep learning models such as GNN models and graph transformer models are neglected, which prompts us to integrate those methods in our benchmark. Also, pretaining strategies are well-known for its ability in utilizing unlabeled data (Lu et al., 2024; Yue et al., 2024).
Challenges. Previous benchmarks related to molecular learning have offered valuable insights regarding their respective libraries and implementation interfaces. DeepPurpose [^1] https://github.com/kexinhuang12345/DeepPurpose (Huang et al., 2020) has provided an interface that implements the task with a majority of drug discovery tasks, which only has protein-protein interaction and protein function prediction implemented. Datasets on proteins are lacking as well. TorchProtein [^2] https://github.com/DeepGraphLearning/PEER_Benchmark  (Xu et al., 2022), also named as PEER, implemented most of the tasks in the protein field. In terms of models, the focus has largely been on sequential learning methods: Convolutional Neural Networks (CNNs), Transformers, and ESM architectures. This suggests that there are still many structural methods available for consideration. Furthermore, PEER’s interface is not user-friendly without prior domain knowledge in graphs and biochemistry. This situation presents an opportunity to improve the existing interface with regard to simplicity and comprehensibility.
Contribution. The major contributions of our paper are summarized as
  • Comprehensive Benchmarking: We curate a benchmark to evaluate the performance of 8 neural network architectures (including CNNs, RNNs, transformers, and various graph neural networks) on these 7 essential protein learning tasks (including protein function prediction, protein localization prediction, protein-protein interaction prediction, antigen epitope prediction, antibody paratope prediction, CRISPR repair outcome prediction, and antibody developability prediciton), demonstrating superior performance and scalability.
  • User-friendly Library: We develop DeepProtein, a specialized deep learning library that integrates these neural network architectures (CNNs, RNNs, transformers, and graph neural networks) on these 7 protein-related tasks such as protein function prediction, protein localization prediction, protein-protein interaction prediction, antigen epitope prediction, and antibody developability prediciton. For each task, our library enables one command line to run every method.
  • Enhanced Accessibility: We provide extensive documentation and tutorials to facilitate user engagement and promote reproducible research, building upon the foundation of the well-known drug discovery library, DeepPurpose (Huang et al., 2020).
    Figure 1: DeepProtein framework.
2 Related Works
@@
user-friendly $\times$ $\times$ $\times$
@@ 3.1 AI-solvable Protein Problems
    • PPI Affinity (Moal and Fernández-Recio, 2012). It consists of 2,682 protein-protein pairs with real-valued groundtruth.
@@
    • SAbDab-Liberis (Liberis et al., 2018) is curated from SAbDab (Dunbar et al., 2014). It consists of 1,023 antibody chain sequences; each antibody contains both heavy and light chain sequences. It can be downloaded from TDC (https://tdcommons.ai/single_pred_tasks/paratope/#sabdab-liberis-et-al).
@@
In this library, we follow the train-validation-test split in PEER benchmark (Xu et al., 2022) and TDC (Huang et al., 2022). Each individual split is reported from Table 2 to 6.
@@ 3.2 Cutting-edge Deep Learning Methods
At the core of deep learning lies the artificial neural network, a machine learning technique inspired by the architecture and functionality of the human brain. What distinguishes deep learning from other machine learning approaches is its exceptional capacity to recognize and analyze complex, nonlinear patterns in data, leading to enhanced performance and accuracy. Concretely, we incorporate several cutting-edge neural network architectures into two groups: 1) sequential-based learning and 2) structural-based learning. Detailed model architectures are described as follows:
Sequential based learning
It generally takes a sequence as an input, uses one-hot encoding to pre-encode the input characters. Such learning methods include convolutional neural network, recurrent neural network and transformers.
  • Convolutional Neural Network (CNN) (One-dimensional) captures the local patterns in the data features, commonly used to analyze images and text. (One-dimensional) Convolutional neural network (CNN) takes amino acid sequences as the input. CNN has four layers; the number of filters for the four layers is 32, 64, and 96 respectively. The kernel sizes are 4, 8, and 12, respectively. The convolutional layer is followed by a one-layer MLP (multi-layer perceptron) to produce the prediction, which is a scalar.
  • Recurrent Neural Network (RNN) models sequence data and captures the long-term dependencies in the sequence data. RNN has two well-known variants: long short-term memory networks (LSTMs) (Hochreiter and Schmidhuber, 1996) and gated recurrent units (GRU) (Cho et al., 2014). The difference between GRU and LSTM is that GRU simplifies LSTM by removing the cell state and reducing the number of gates. We use a two-layer bi-directional GRU following three-layer CNN as the neural network architecture. The dimension of the hidden state in GRU is set to 64. ReLU function is applied after each GRU or CNN layer.
@@
    Structural-based learning
    It generally transforms the input sequence into a valid SMILES string, then transforms the chemical substance into a graph. Then, graph filters are learned toward the input graph signal. Such learning methods are widely called Graph Neural Networks. Recently, graph transformers have shown their power in protein function prediction, and we included them as a part of structural-based learning.
@@
    • Graph Convolutional Network (GCN) (Kipf and Welling, 2016). GCN is a GNN variant that iteratively updates the node representation by aggregating the information from its neighbors. GCN has three layers, and the node embedding dimension is set to 64. After GCN, all the node embeddings are aggregated with a readout function (Weighted Sum and Max) to get graph-level embedding, followed by a one-layer MLP to get the final prediction. BatchNorm is applied after MLP layers.
@@
    • Neural Fingerprint (NeuralFP) (Duvenaud et al., 2015). NeuralFP uses Graph convolutional network (GCN) (Kipf and Welling, 2016) to learn a neural network-based molecular embedding (also known as molecular neural fingerprint, or NeuralFP) from a large amount of molecule data without labels. The neural fingerprint is essentially a real-valued vector, also known as embedding. Then, the neural fingerprint is fixed and fed into a three-layer MLP to make the prediction. Node embedding dimension is 64. BatchNorm is applied after MLP layers.
@@
Training setup.
For all the models, the maximal training epoch number is set to 100. We employed the Adam optimizer (Kingma and Ba, 2014) for training, with a default learning rate of 0.0001 for sequence-based learning and 0.00001 for structural-based learning. The batch size is equal to 32. More detailed hyper-parameter setups are listed in Table 7 in the appendix.

### v2

@@ Abstract
Deep learning has deeply influenced protein science, enabling breakthroughs in predicting protein properties, higher-order structures, and molecular interactions. This paper introduces DeepProtein, a comprehensive and user-friendly deep learning library tailored for protein-related tasks. It enables researchers to seamlessly address protein data with cutting-edge deep learning models. To assess model performance, we establish a benchmark evaluating different deep learning architectures across multiple protein-related tasks, including protein function prediction, subcellular localization prediction, protein-protein interaction prediction, and protein structure prediction. Furthermore, we introduce DeepProt-T5, a series of fine-tuned Prot-T5-based models that achieve state-of-the-art performance on four benchmark tasks, while demonstrating competitive results on six of others. Comprehensive documentation and tutorials are available which could ensure accessibility and support reproducibility. Built upon the widely used drug discovery library DeepPurpose, DeepProtein is publicly available at https://github.com/jiaqingxie/DeepProtein.
@@ 1 Introduction
Understanding the representation of proteomics is vital in developing traditional biological and medical progress (Wu et al., 2022b; Fu et al., 2024), multi-omics genomics (Wu et al., 2022a; Chen et al., 2021), and curing human diseases (Chen et al., 2024b; c). Being the working house of the cell, it provides many functions that support human daily life, such as catalyzing biochemical reactions that occur in the body as a role of enzymes and providing helpful immune responses against harmful substances that act as immunoglobulin (Wu et al., 2024). Under the necessity of analyzing those useful proteins, several related protein databases are available to researchers (Berman et al., 2000; Bairoch & Apweiler, 2000; Consortium, 2015; Pontén et al., 2008). Apart from the 2D database, some recent 3D Protein Database used AlphaFold 2.0 (Jumper et al., 2021) is important to better assist in learning those representations in 3d-dimensional space. The success of AlphaFold 2.0 has sparked a significant increase in interest in using machine learning techniques for protein learning tasks, of which the goal is to improve our understanding of proteins’ biochemical mechanisms.
Deep learning has revolutionized protein science, driving significant advancements in various protein-related tasks. These include protein-protein interactions (Gainza et al., 2020), protein folding (Jumper et al., 2021; Lu, 2022; Panou & Reczko, 2020; Chen et al., 2016), protein-ligand interactions (Li et al., 2021b; Xia et al., 2023), and protein function and property prediction (Gligorijević et al., 2021; Sevgen et al., 2023). The development of deep neural architectures has played a crucial role in these tasks, with approaches leveraging both sequence-based and structure-based models. Sequence-based models, such as convolutional neural networks (CNNs) (Shanehsazzadeh et al., 2020) and transformers, have shown strong performance in protein learning tasks. The TAPE Transformer (Rao et al., 2019) and pre-trained transformer models such as ProtBERT (Brandes et al., 2022) have demonstrated the effectiveness of self-supervised learning in capturing protein sequence representations. Beyond sequence-based methods, structure-based deep learning has gained traction with graph neural networks (GNNs), which leverage 3D structural information to enhance structural property (Jing et al., 2020; Zhang et al., 2022). Recently, graph transformers have emerged as a powerful alternative, combining the advantages of transformers (global attention) and message-passing neural networks (sparse attention) to model protein structures more effectively (Yuan et al., 2022; Gu et al., 2023).
While transformers have been considered state-of-the-art in previous benchmarks (Xu et al., 2022), comprehensive comparisons between CNN, transformer, GNN, and other advanced architectures remain under-explored. This gap motivates us to systematically integrate and evaluate these methods in our benchmark. Furthermore, pretraining strategies have been prevailing in protein science, which have utilized the large-scale unlabeled protein data to improve downstream performance (Lu et al., 2024; Yue et al., 2024). With the advent of large foundation models, protein properties can now be inferred through prompt engineering, such as BioMistral (Labrak et al., 2024), BioT5/BioT5+ (Pei et al., 2023; 2024), and ChemLLM (Zhang et al., 2024b). Both advancements in molecule pretraining and question-answering language models in molecules brought more possibilities in the field of protein engineering.
@@
Fold $\times$ $\times$ $\times$ $\times$
Secondary Structure $\times$ $\times$ $\times$
Challenges. Previous benchmarks related to molecular learning have offered valuable insights regarding their respective libraries and implementation interfaces. DeepPurpose [^1] https://github.com/kexinhuang12345/DeepPurpose (Huang et al., 2020) has provided an interface that implements the task with a majority of drug discovery tasks, which only has protein-protein interaction and protein function prediction implemented. Datasets on proteins are lacking as well. TorchProtein [^2] https://github.com/DeepGraphLearning/PEER_Benchmark  (Xu et al., 2022), also named as PEER, implemented most of the tasks in the protein field. In terms of models, the focus has largely been on sequence-based methods: Convolutional Neural Networks (CNNs), Transformers, and ESM architectures. This suggests that there are still many structure-based methods (GNN) or pre-trained protein language models available (such as ProtBert or Prot-T5) for consideration. Furthermore, PEER’s interface is not user-friendly without prior domain knowledge in graphs and biochemistry. This presents an opportunity to improve the existing interface regarding simplicity and comprehensibility.
Contribution. Our key contributions are summarized as:
  • Comprehensive Benchmarking: We curate a benchmark to evaluate the performance of eight coarse grained deep learning architectures, including CNNs, CNN-RNNs, RNNs, transformers, graph neural networks, graph transformers, pre-trained protein language models, and large language models. This benchmark covers eight protein learning tasks, including protein function prediction, protein localization prediction, protein-protein interaction prediction, antigen epitope prediction, antibody paratope prediction, CRISPR repair outcome prediction, antibody developability prediction, and protein structure prediction. Our benchmark demonstrates the strength, scalability and limitation of the mentioned approach, respectively.
  • User-friendly Library: We develop DeepProtein, a specialized deep learning library that integrates these neural network architectures for protein-related tasks. DeepProtein offers a simple, command-line interface for running models on all supported tasks, making it accessible to researchers with minimal deep learning expertise.
  • Enhanced Accessibility: We provide comprehensive documentation, tutorials, and pre-configured pipelines. Inherited from DeepPurpose (Huang et al., 2020), our library ensures seamless integration with existing protein frameworks or personalized protein databases, and enables reproducibility.
  • Fine-Tuned Models – DeepProt-T5: We have released our fine-tuned Prot-T5-XL models for each task, which is available on HuggingFace. The model family is called DeepProt-T5. These models achieve either state-of-the-art or competitive performance across our DeepProtein benchmark, so there is no need for the redundant retraining process, making model deployment much more efficient and convenient.
2 Related Works
Figure 1: DeepProtein framework. Part 1. A. DeepProtein is mainly selected from TorchDrug and Therapeutics Data Commons (TDC), where only protein tasks are considered, and all drug-related tasks are excluded, such as drug-target interactions. Specifically, DeepPurpose has established such a pipeline in their library. B. Both sequence-based and structure-based methods are included in DeepProtein. For some graph neural networks, we utilized edge featurizers to generate additional edge information since inputs are 2-dimensional. Protein language models, large language models, and our pre-trained T5 (DeepProt-T5) are discussed in Figure 2. C. Task types are: protein function prediction, subcellular localization prediction, protein-protein interaction prediction, and protein structure prediction. They can be classified as either a 1 (protein)-to-1 (aim) problem or a 2 (proteins)-to-1 (aim) problem, which meets the researchers’ needs. D. An earlier version of DeepProtein could be executed within 20 lines of code. The newest version of DeepProtein could be executed within 10 lines of code, where we further wrapped the data processing steps. Researchers can also provide their own data to either train or perform inference with the help of DeepProtein. E. In this paper, we provide comprehensive results, including the performance of each model on corresponding tasks, the differences among sequence-based models, structure-based models, and pre-trained protein language models, and the computation resources, including time-stamps and GPU memory assumptions. As DeepProtein supports wandb, we also provide two wandb repositories that record the results of all experiments, which are https://wandb.ai/jiaqing/DeepProtein?nw=nwuserjiaqing and https://wandb.ai/jiaqing/DeepPurposePP. Tables and figures are presented later in this paper.
@@ 3.1 AI-solvable Protein Problems
    • PPI Affinity (Moal & Fernández-Recio, 2012). It consists of 2,682 protein-protein pairs with real-valued groundtruth.
@@
  • Protein Structure Prediction. Protein structure prediction (PSP) is a fundamental problem in computational biology, aiming to determine the three-dimensional structure of a protein from its amino acid sequence. Specifically, in our benchmark, the task is to predict the family of a folding or secondary structure family that it belongs to. Since a protein’s structure dictates its function, accurate prediction is crucial for understanding the topology of the protein. PSP can be broadly divided into global topology prediction and local structural prediction, which include tasks such as fold classification and secondary structure prediction:
    • Fold (Hou et al., 2018). This task involves predicting the global structural topology of a protein, categorizing it into one of the predefined fold classes (within 0, 1, …, 1194). During inference, we predict the class of each protein’s superfamily. It contains 13766 samples overall.
    • Secondary Structure (Klausen et al., 2019). This task focuses on predicting the local structural elements (coil, strand, helix) of each residue in a protein sequence. It serves as an intermediate step for more complex structure prediction tasks and is useful in applications such as functional analysis and multiple sequence alignment. This is a residue-level 3-class classification problem where the number of samples is equal to 11361.
In this library, we follow the train-validation-test split in PEER benchmark (Xu et al., 2022) and TDC (Huang et al., 2022). Each individual split is reported from Table 2 to 7.
Figure 2: DeepProtein framework, and DeepProt-T5. Part 2. A. For pre-trained protein language models, we directly use the initial string as the input instead of the transformed SMILES , since most of the protein language models have learned such representations. Strings are tokenized and carried as inputs to the model. We fine-tuned the models on the downstream tasks with the fixed embeddings, which means that after feature extraction, the upstream model parameters will not be trained any more. B. For large language models, such as ChemLLM and LlaSMol, we did not fine-tune them on the downstream tasks due to limited GPU resources. Instead we directly performed inference on downstream tasks where appropriate prompt templates should be carefully designed here. C. We extend the family of Prot-T5 to our DeepProt-T5, where the upstream embeddings are dynamic, by using the huggingface trainer. All fine-tuned models could be found here:https://huggingface.co/jiaxie.
@@ 3.2 Cutting-edge Deep Learning Methods
At the core of deep learning lies the artificial neural network, a machine learning technique inspired by the architecture and functionality of the human brain. What distinguishes deep learning from other machine learning approaches is its exceptional ability to recognize and analyze complex, non-linear patterns in data, leading to enhanced performance and accuracy. Concretely, we incorporate several cutting-edge neural network architectures into two groups: 1) sequential-based learning and 2) structural-based learning. Detailed model architectures are described as follows:
Sequential based learning It generally takes a sequence as an input and uses one-hot encoding to pre-encode the input characters. Such learning methods include convolutional neural networks, recurrent neural networks, and transformers.
  • Convolutional Neural Network (CNN) (One-dimensional) captures the local patterns in the data features, commonly used to analyze images and text. (One-dimensional) Convolutional neural network (CNN) takes amino acid sequences as the input. CNN has four layers; the number of filters for the four layers is 32, 64, and 96, respectively. The kernel sizes are 4, 8, and 12, respectively. The convolutional layer is followed by a one-layer MLP (multi-layer perceptron) to predict as a scalar.
  • Recurrent Neural Network (RNN) models sequence data and captures the long-term dependencies in the sequence data. RNN has two well-known variants: long short-term memory networks (LSTMs) (Hochreiter & Schmidhuber, 1996) and gated recurrent units (GRU) (Cho et al., 2014). The difference between GRU and LSTM is that GRU simplifies LSTM by removing the cell state and reducing the number of gates. We use a two-layer bi-directional GRU following three-layer CNN as the neural network architecture. The dimension of the hidden state in GRU is set to 64. ReLU function is applied after each GRU or CNN layer.
@@
    Structural-based learning It generally transforms the input sequence into a valid SMILES string, then transforms the chemical substance into a graph. Then, graph filters are learned toward the input graph signal. Such learning methods are widely called Graph Neural Networks. Recently, graph transformers have shown their power in protein function prediction, and we included them as a part of structural-based learning.
@@
    • Graph Convolutional Network (GCN) (Kipf & Welling, 2016). GCN is a GNN variant that iteratively updates the node representation by aggregating the information from its neighbors. GCN has three layers, and the node embedding dimension is set to 64. After GCN, all the node embeddings are aggregated with a readout function (Weighted Sum and Max) to get graph-level embedding, followed by a one-layer MLP to get the final prediction. BatchNorm is applied after MLP layers.
@@
    • Neural Fingerprint (NeuralFP) (Duvenaud et al., 2015). NeuralFP uses Graph convolutional network (GCN) (Kipf & Welling, 2016) to learn a neural network-based molecular embedding (also known as molecular neural fingerprint, or NeuralFP) from a large amount of molecule data without labels. The neural fingerprint is essentially a real-valued vector, also known as embedding. Then, the neural fingerprint is fixed and fed into a three-layer MLP to make the prediction. Node embedding dimension is 64. BatchNorm is applied after MLP layers.
@@
  • Foundation Model. A foundation model is a large-scale, pre-trained machine learning model trained on extensive and diverse datasets, typically using self-supervised or unsupervised learning techniques. These models learn generalizable features and patterns from data, allowing them to perform various downstream tasks with minimal task-specific fine-tuning.
    • ESM. The Evolutionary Scale Modeling (ESM) utilizes large-scale pretraining on vast protein sequence data to capture evolutionary relationships and functional patterns within proteins (Lin et al., 2023; Rives et al., 2021). It benefits from Masked Language Modeling (MLM) and the transformer architecture. In this paper, we consider two ESM variants with different model sizes: ESM-1b and ESM-2-650M. The latter incorporates Rotary Position Embedding (RoPE) within the ESM-1 framework. We evaluate both models, where the embedding size is 1280.
    • Prot-T5-XL. First introduced in ProtTrans (Elnaggar et al., 2021), Prot-T5-XL-UniRef50 is based on the T5-3B model and was pre-trained on a large corpus of protein sequences using a self-supervised approach. A key difference from the original T5 model is the denoising objective: while the original T5-3B model used a span denoising objective, this model employs a BART-like MLM denoising objective. The masking probability follows the original T5 training, randomly masking 15% of the amino acids in the input. The embedding dimension is 1024.
    • ProtBert. In addition to Prot-T5, ProtTrans includes a model pre-trained on BERT. Bidirectional Encoder Representations from Transformers (BERT) is a transformer-based neural network architecture pre-trained on unlabeled sequence data (Devlin et al., 2019). A key difference between ProtBert and the original BERT is how sequences are treated as separate documents, eliminating the need for next sentence prediction. The masking strategy follows the original BERT training, where 15% of the amino acids in the input are randomly masked. The embedding dimension is 1024.
  • Large Language Model. In this paper, we distinguish foundation models as a class of pre-trained protein language models, whereas we define large language models (LLMs) as decoder-only models designed to generate sequential responses regarding the properties of one or multiple proteins given their sequences and a specific prompt. Due to computational constraints, fine-tuning 7B-scale models is resource-intensive; therefore, we focus on evaluating their performance instead. We consider two Protein LLMs in our study: ChemLLM-7B and LlaSMol-Mistral-7B. Their generalization ability to protein-related tasks has not been explored before. Additionally we provide the chat and prompt template in the appendix.
    • ChemLLM-7B. The backbone of ChemLLM-7B (Zhang et al., 2024a) is the InternLM2-Base-7B model. It was initially trained on the Multi-Corpus-1.7M dataset on Hugging Face and later fine-tuned using instruction-tuning methods on ChemData (7M) and Multi-Corpus (1.7M). ChemLLM-7B has demonstrated superior performance over GPT-4 in retrosynthesis and temperature prediction tasks. The model provides a set of predefined instruction-following templates, which are used in our study and detailed in the appendix.
    • LlaSMol-Mistral-7B. The backbone of LlaSMol-Mistral-7B (Yu et al., 2024) is Mistral-7B. It was fine-tuned using SMolInstruct, a large-scale, high-quality dataset designed for instruction tuning. SMolInstruct comprises 14 selected chemistry-related tasks and over three million samples, providing a solid foundation for training and evaluating LLMs in the field of chemistry. Specifically, in our experiments, we wrap the input SMILES sequences with $\rangle\langle\text{/PROTEIN}\rangle$PROTEIN ⟩ token pairs to adapt them for protein-related tasks. The template is detailed in the appendix.
  • DeepProt-T5. We have trained Prot-T5-XL on our benchmark datasets individually (Figure 2). Different from the fixed embeedings, dynamic embeddings enabled us to finetuned the upstream architectures while also maintain a good predictive power for downstream tasks. This leads to a series of DeepProt-T5 models.
Training setup. For all models, the maximal training epoch number is set to 100. We used Adam optimizer (Kingma & Ba, 2014) for training, with a default learning rate of 0.0001 for sequence-based learning and 0.00001 for structural-based learning. The batch size is equal to 32. More detailed hyper-parameter setups are listed in Table 10 in the appendix. For DeepProt-T5 models, fine-tuning has used a more generalized learning rate of 0.00002 for all models, batch size is equal to 10 to avoid memory errors.
Figure 3: Results of two metrics for selected deep learning methods for DeepProtein Benchmark. For regression task, metrics are Spearman (Pearson) Coefficient and Mean Absolute Error (MAE). For the binary classification task, metrics are ROC (or PR-AUC) and averaged macro F1. For multi classification task, metrics are the accuracy and averaged macro F1. Our DeepProt-T5 are competitive among deep learning methods included in our benchmark, and have improved original Prot-T5 models on six tasks: Beta-lactamase, Solubility, SubCellular, PPI_Affinity, CRISPR and Fold.